# FMCW Radar Principles and Human Activity Recognition Systems: Foundations, Techniques, and Applications

### v1

@@ 1.1 Why Python for FMCW Radar?
Python has become one of the most popular programming languages for scientific and engineering applications [61], including radar signal processing. Its strengths, such as ease of use, an extensive collection of libraries, and a strong community of users, make it an ideal choice for FMCW (Frequency Modulated Continuous Wave) radar development [35].
@@ 1.2 Setting Up the Python Environment
Setting up a Python environment suitable for FMCW radar signal processing is crucial for ensuring a smooth development process. The recommended way to set up your environment is by using Anaconda, a powerful distribution that simplifies package management and deployment [9]. This section will guide you through installing the necessary tools and libraries for FMCW radar development.
@@ 3.1 Introduction to NumPy Arrays
NumPy is a fundamental package for scientific computing [36] in Python and plays a crucial role in radar signal processing. Its core feature is the powerful N-dimensional array object, known as the ndarray, which is efficient for handling large-scale, multi-dimensional radar data.
@@ 4.1 Introduction to Matplotlib
Matplotlib is a powerful plotting library [55] in Python that allows users to create various types of static, interactive, and animated visualizations. For radar signal processing, visualization is an essential part of analyzing and interpreting signals at different stages, such as in the time domain and frequency domain. This section introduces Matplotlib and demonstrates how to use it for radar signal visualization.
@@ 5.1 Classes and Objects in Python
In Python, a class is a blueprint for creating objects [23]. An object is an instance of a class, containing both attributes (data) and methods (functions) that operate on the data. Classes allow you to encapsulate radar-related functionalities into logical units. For example, you might have a class that represents a radar sensor, a signal processor, or even a specific radar waveform.
@@ 5.3 When and How to Use OOP for Radar Applications
However, OOP might not be necessary for simpler radar applications where the focus is on single functions or scripts. In such cases, functional programming [25], where you simply define functions and call them, might be more appropriate. You should avoid unnecessary complexity when a function-based approach is sufficient for tasks like signal filtering or generating radar waveforms.
@@ 6.4 Asynchronous Programming
Asynchronous programming [21] is a programming paradigm that allows you to run tasks in a non-blocking manner, meaning that tasks can execute concurrently without waiting for each other to complete. In Python, the asyncio library provides a framework for writing asynchronous code.
@@ 7.1 Why MATLAB for FMCW Radar?
MATLAB (short for MATrix LABoratory) is a powerful programming environment [40] designed for numerical computing and visualization. Its extensive library of built-in functions and toolboxes makes it particularly useful for engineers working in fields such as signal processing, control systems, and radar systems. When applied to FMCW (Frequency Modulated Continuous Wave) radar, MATLAB provides several distinct advantages for both beginners and experienced engineers alike.
@@ 9.1 Time-Domain Signal Analysis
Time-domain signal analysis focuses on analyzing signals [38] as they evolve over time. For FMCW (Frequency Modulated Continuous Wave) radar systems, the time-domain representation of signals is fundamental to understanding how the radar transmits and receives waveforms. MATLAB offers powerful functions for working with time-domain signals, and this section introduces some of the most important tools and techniques.
@@ 1. Signal Smoothing
Signal smoothing [16] is a common operation that removes high-frequency noise from the signal, making the underlying pattern easier to identify. In MATLAB, this can be achieved using moving average filters or other smoothing techniques. Below is an example of how to smooth a noisy signal using a moving average filter:
@@ 9.2.1 Fourier Transform with MATLAB
The Fast Fourier Transform [39] (FFT) is a computational tool that converts a time-domain signal into its frequency-domain representation. For radar systems, the FFT helps in analyzing how different frequency components contribute to the overall signal.
@@ 9.2.2 Power Spectral Density (PSD) Estimation
Power Spectral Density (PSD) provides information [10] about how the signal’s power is distributed across different frequencies. In radar systems, PSD is useful for identifying signal components and background noise levels.
@@ 9.3.1 Designing Filters in MATLAB
MATLAB provides several ways to design filters, including low-pass, high-pass, and band-pass filters, which are useful for isolating specific frequency components of radar signals. The ‘designfilt‘ [14] function can be used for this purpose.
@@ 10.1 Simulating FMCW Waveforms
Frequency Modulated Continuous Wave [68] (FMCW) radar transmits signals whose frequency increases linearly over time during a specific duration, known as the sweep period. The key parameters for generating FMCW waveforms include:
@@ 10.2.2 Velocity Estimation via Doppler Effect
The Doppler effect [37]causes a frequency shift in the received signal due to the relative motion between the radar and the target. For an FMCW radar, the Doppler frequency shift $f_{D}$ is related to the velocity of the target:
@@ 11.1 Parallel Computing in MATLAB
When dealing with large-scale radar data, processing can become computationally expensive. MATLAB provides powerful tools for parallel computing [53]that allow users to take advantage of multi-core processors, GPUs, and distributed computing environments to speed up data processing tasks.
@@ 12.1 Why and How Use Deep Learning for Radar Signal Processing?
Radar systems, particularly Frequency Modulated Continuous Wave (FMCW) radars, are essential in various applications such as automotive safety, aerospace, maritime navigation, and surveillance. Traditional radar signal processing techniques have served these applications well, but they have limitations when dealing with complex environments and large volumes of data. Deep learning offers powerful tools to overcome these challenges by automatically extracting features and learning intricate patterns from data [30].
@@ 12.1.5 Example: Using a CNN for Target Classification
4class RadarCNN(nn.Module):
@@
7 self.conv1 = nn.Conv2d(1, 16, 3) # Input channels, output channels, kernel size
8 self.conv2 = nn.Conv2d(16, 32, 3)
9 self.fc1 = nn.Linear(32 * 6 * 6, 128) # Adjust dimensions accordingly
10 self.fc2 = nn.Linear(128, num_classes)
@@ 12.1.6 Example: Human Activity Recognition with LSTM Networks
1class RadarLSTM(nn.Module):
@@
4 self.lstm = nn.LSTM(input_size, hidden_size, num_layers, batch_first=True)
5 self.fc = nn.Linear(hidden_size, num_classes)
@@ 12.1.7 Advanced Techniques: Transformers and PointNet
3class RadarTransformer(nn.Module):
@@
7 self.fc = nn.Linear(input_size, num_classes)
@@
1class PointNet(nn.Module):
@@
4 self.fc1 = nn.Linear(3, 64)
5 self.fc2 = nn.Linear(64, 128)
6 self.fc3 = nn.Linear(128, 1024)
7 self.fc4 = nn.Linear(1024, num_classes)
@@ 13.1.1 Convolutional Layers
4class BasicCNN(nn.Module):
@@
7 self.conv1 = nn.Conv2d(1, 16, kernel_size=3, stride=1, padding=1)
8 self.conv2 = nn.Conv2d(16, 32, kernel_size=3, stride=1, padding=1)
9 self.fc1 = nn.Linear(32 * 8 * 8, 128)
10 self.fc2 = nn.Linear(128, 2) # Binary classification
@@
13 x = nn.functional.relu(self.conv1(x))
14 x = nn.functional.max_pool2d(x, 2)
15 x = nn.functional.relu(self.conv2(x))
16 x = nn.functional.max_pool2d(x, 2)
@@
18 x = nn.functional.relu(self.fc1(x))
@@ 13.3.1 Transfer Learning for Radar Applications
11resnet.fc = nn.Linear(resnet.fc.in_features, num_classes)
@@ 14.1.2 LSTM for Radar Signal Processing
4class RadarLSTM(nn.Module):
@@
7 self.lstm = nn.LSTM(input_size, hidden_size, num_layers, batch_first=True)
8 self.fc = nn.Linear(hidden_size, num_classes)
@@ 14.2.1 CNN+LSTM Model for Human Activity Recognition (HAR)
1class CNN_LSTM(nn.Module):
@@
4 self.conv1 = nn.Conv2d(1, 16, kernel_size=3, stride=1, padding=1)
5 self.conv2 = nn.Conv2d(16, 32, kernel_size=3, stride=1, padding=1)
6 self.lstm = nn.LSTM(32 * 8 * 8, hidden_size, num_layers, batch_first=True)
7 self.fc = nn.Linear(hidden_size, num_classes)
@@
12 x = nn.functional.relu(self.conv1(x))
13 x = nn.functional.max_pool2d(x, 2)
14 x = nn.functional.relu(self.conv2(x))
15 x = nn.functional.max_pool2d(x, 2)
@@ 14.2.4 Conclusion
Radar systems, particularly modern FMCW radars, generate point cloud data that captures 3D spatial information. Point clouds are sets of points in a 3D space that represent the surfaces and objects detected by the radar. Unlike traditional image-based data, point clouds are unordered and irregular, making them challenging to process using conventional deep learning models. PointNet [45] and PointNet++ are two architectures specifically designed to handle point cloud data, making them well-suited for tasks such as object detection, segmentation, and classification in radar systems.
@@ 15.1.1 PointNet Architecture
5class PointNet(nn.Module):
@@
8 self.fc1 = nn.Linear(3, 64)
9 self.fc2 = nn.Linear(64, 128)
10 self.fc3 = nn.Linear(128, 1024)
11 self.fc4 = nn.Linear(1024, 512)
12 self.fc5 = nn.Linear(512, 256)
13 self.fc6 = nn.Linear(256, num_classes)
@@ 15.2.1 Hierarchical Learning in PointNet++
1class PointNetPlusPlus(nn.Module):
@@
4 self.fc1 = nn.Linear(3, 64)
5 self.fc2 = nn.Linear(64, 128)
6 self.fc3 = nn.Linear(128, 1024)
7 self.fc4 = nn.Linear(1024, 512)
8 self.fc5 = nn.Linear(512, 256)
9 self.fc6 = nn.Linear(256, num_classes)
@@ 16.1.1 Self-Attention Mechanism
At the core of the Transformer architecture is the self-attention mechanism [70], which allows the model to examine the entire input sequence simultaneously and assign different attention weights to each element in the sequence. This enables the model to focus on important parts of the data, regardless of their position in the sequence.
@@
5class SelfAttention(nn.Module):
@@
10 self.values = nn.Linear(embed_size, embed_size, bias=False)
11 self.keys = nn.Linear(embed_size, embed_size, bias=False)
12 self.queries = nn.Linear(embed_size, embed_size, bias=False)
13 self.fc_out = nn.Linear(embed_size, embed_size)
@@ 16.2.1 Transformer-based Models for Radar Data
1class TransformerRadarModel(nn.Module):
@@
4 self.self_attention_layers = nn.ModuleList([
@@ 5 SelfAttention(embed_size, heads) for _ in range(num_layers)
7 self.fc = nn.Linear(embed_size, num_classes)
@@ 16.3.1 Self-Attention in 3D Point Cloud Data
5class TransformerPointCloud(nn.Module):
@@
8 self.attention = nn.MultiheadAttention(embed_dim=embed_size, num_heads=heads)
9 self.fc1 = nn.Linear(embed_size, 512)
10 self.fc2 = nn.Linear(512, num_classes)
@@ 16.4.2 Transformers for Long-Range Dependencies
1class PointNetPlusPlusTransformer(nn.Module):
@@
5 self.attention = nn.MultiheadAttention(embed_dim=embed_size, num_heads=heads)
6 self.fc = nn.Linear(embed_size, num_classes)
@@ 17.1.2 ResNet for Feature Extraction
ResNet (Residual Networks) is one of the most popular pre-trained models for feature extraction [2]. It introduces skip connections, allowing the network to bypass certain layers and avoid the vanishing gradient problem. This architecture enables the model to be very deep, making it effective for capturing intricate patterns in radar data.
@@
9resnet = nn.Sequential(*list(resnet.children())[:-1])

### v2

@@
   Xinyuan Song
Emory University
xsong30@emory.edu
@@ 1.1 Why Python for FMCW Radar?
Python has become one of the most popular programming languages for scientific and engineering applications [van2007python], including radar signal processing. Its strengths, such as ease of use, an extensive collection of libraries, and a strong community of users, make it an ideal choice for FMCW (Frequency Modulated Continuous Wave) radar development [marshall2008fmcw].
@@ 1.2 Setting Up the Python Environment
Setting up a Python environment suitable for FMCW radar signal processing is crucial for ensuring a smooth development process. The recommended way to set up your environment is by using Anaconda, a powerful distribution that simplifies package management and deployment [clark2022standardizing]. This section will guide you through installing the necessary tools and libraries for FMCW radar development.
@@ 3.1 Introduction to NumPy Arrays
NumPy is a fundamental package for scientific computing [mehta2015mastering] in Python and plays a crucial role in radar signal processing. Its core feature is the powerful N-dimensional array object, known as the ndarray, which is efficient for handling large-scale, multi-dimensional radar data.
@@ 4.1 Introduction to Matplotlib
Matplotlib is a powerful plotting library [sial2021comparative] in Python that allows users to create various types of static, interactive, and animated visualizations. For radar signal processing, visualization is an essential part of analyzing and interpreting signals at different stages, such as in the time domain and frequency domain. This section introduces Matplotlib and demonstrates how to use it for radar signal visualization.
@@ 5.1 Classes and Objects in Python
In Python, a class is a blueprint for creating objects [hernandez2021python]. An object is an instance of a class, containing both attributes (data) and methods (functions) that operate on the data. Classes allow you to encapsulate radar-related functionalities into logical units. For example, you might have a class that represents a radar sensor, a signal processor, or even a specific radar waveform.
@@ 5.3 When and How to Use OOP for Radar Applications
However, OOP might not be necessary for simpler radar applications where the focus is on single functions or scripts. In such cases, functional programming [hughes1989functional], where you simply define functions and call them, might be more appropriate. You should avoid unnecessary complexity when a function-based approach is sufficient for tasks like signal filtering or generating radar waveforms.
@@ 6.4 Asynchronous Programming
Asynchronous programming [hattingh2020using] is a programming paradigm that allows you to run tasks in a non-blocking manner, meaning that tasks can execute concurrently without waiting for each other to complete. In Python, the asyncio library provides a framework for writing asynchronous code.
@@ 7.1 Why MATLAB for FMCW Radar?
MATLAB (short for MATrix LABoratory) is a powerful programming environment [otto2005introduction] designed for numerical computing and visualization. Its extensive library of built-in functions and toolboxes makes it particularly useful for engineers working in fields such as signal processing, control systems, and radar systems. When applied to FMCW (Frequency Modulated Continuous Wave) radar, MATLAB provides several distinct advantages for both beginners and experienced engineers alike.
@@ 9.1 Time-Domain Signal Analysis
Time-domain signal analysis focuses on analyzing signals [nguyen2014time] as they evolve over time. For FMCW (Frequency Modulated Continuous Wave) radar systems, the time-domain representation of signals is fundamental to understanding how the radar transmits and receives waveforms. MATLAB offers powerful functions for working with time-domain signals, and this section introduces some of the most important tools and techniques.
@@ 1. Signal Smoothing
Signal smoothing [enke1976signal] is a common operation that removes high-frequency noise from the signal, making the underlying pattern easier to identify. In MATLAB, this can be achieved using moving average filters or other smoothing techniques. Below is an example of how to smooth a noisy signal using a moving average filter:
@@ 9.2.1 Fourier Transform with MATLAB
The Fast Fourier Transform [nussbaumer1982fast] (FFT) is a computational tool that converts a time-domain signal into its frequency-domain representation. For radar systems, the FFT helps in analyzing how different frequency components contribute to the overall signal.
@@ 9.2.2 Power Spectral Density (PSD) Estimation
Power Spectral Density (PSD) provides information [czifra2009sensitivity] about how the signal’s power is distributed across different frequencies. In radar systems, PSD is useful for identifying signal components and background noise levels.
@@ 9.3.1 Designing Filters in MATLAB
MATLAB provides several ways to design filters, including low-pass, high-pass, and band-pass filters, which are useful for isolating specific frequency components of radar signals. The ‘designfilt‘ [dong2019developing] function can be used for this purpose.
@@ 10.1 Simulating FMCW Waveforms
Frequency Modulated Continuous Wave [zheng2004analysis] (FMCW) radar transmits signals whose frequency increases linearly over time during a specific duration, known as the sweep period. The key parameters for generating FMCW waveforms include:
@@ 10.2.2 Velocity Estimation via Doppler Effect
The Doppler effect [neipp2003analysis]causes a frequency shift in the received signal due to the relative motion between the radar and the target. For an FMCW radar, the Doppler frequency shift $f_{D}$ is related to the velocity of the target:
@@ 11.1 Parallel Computing in MATLAB
When dealing with large-scale radar data, processing can become computationally expensive. MATLAB provides powerful tools for parallel computing [sharma2009matlab]that allow users to take advantage of multi-core processors, GPUs, and distributed computing environments to speed up data processing tasks.
@@ 12.1 Why and How Use Deep Learning for Radar Signal Processing?
Radar systems, particularly Frequency Modulated Continuous Wave (FMCW) radars, are essential in various applications such as automotive safety, aerospace, maritime navigation, and surveillance. Traditional radar signal processing techniques have served these applications well, but they have limitations when dealing with complex environments and large volumes of data. Deep learning offers powerful tools to overcome these challenges by automatically extracting features and learning intricate patterns from data [lang2020comprehensive].
@@ 12.1.5 Example: Using a CNN for Target Classification
4class RadarCNN($n$.Module):
@@
7 self.conv1 = $n$.Conv2d(1, 16, 3) # Input channels, output channels, kernel size
8 self.conv2 = $n$.Conv2d(16, 32, 3)
9 self.fc1 = $n$.Linear(32 * 6 * 6, 128) # Adjust dimensions accordingly
10 self.fc2 = $n$.Linear(128, num_classes)
@@ 12.1.6 Example: Human Activity Recognition with LSTM Networks
1class RadarLSTM($n$.Module):
@@
4 self.lstm = $n$.LSTM(input_size, hidden_size, num_layers, batch_first=True)
5 self.fc = $n$.Linear(hidden_size, num_classes)
@@ 12.1.7 Advanced Techniques: Transformers and PointNet
3class RadarTransformer($n$.Module):
@@
7 self.fc = $n$.Linear(input_size, num_classes)
@@
1class PointNet($n$.Module):
@@
4 self.fc1 = $n$.Linear(3, 64)
5 self.fc2 = $n$.Linear(64, 128)
6 self.fc3 = $n$.Linear(128, 1024)
7 self.fc4 = $n$.Linear(1024, num_classes)
@@ 13.1.1 Convolutional Layers
4class BasicCNN($n$.Module):
@@
7 self.conv1 = $n$.Conv2d(1, 16, kernel_size=3, stride=1, padding=1)
8 self.conv2 = $n$.Conv2d(16, 32, kernel_size=3, stride=1, padding=1)
9 self.fc1 = $n$.Linear(32 * 8 * 8, 128)
10 self.fc2 = $n$.Linear(128, 2) # Binary classification
@@
13 x = $n$.functional.relu(self.conv1(x))
14 x = $n$.functional.max_pool2d(x, 2)
15 x = $n$.functional.relu(self.conv2(x))
16 x = $n$.functional.max_pool2d(x, 2)
@@
18 x = $n$.functional.relu(self.fc1(x))
@@ 13.3.1 Transfer Learning for Radar Applications
11resnet.fc = $n$.Linear(resnet.fc.in_features, num_classes)
@@ 14.1.2 LSTM for Radar Signal Processing
4class RadarLSTM($n$.Module):
@@
7 self.lstm = $n$.LSTM(input_size, hidden_size, num_layers, batch_first=True)
8 self.fc = $n$.Linear(hidden_size, num_classes)
@@ 14.2.1 CNN+LSTM Model for Human Activity Recognition (HAR)
1class CNN_LSTM($n$.Module):
@@
4 self.conv1 = $n$.Conv2d(1, 16, kernel_size=3, stride=1, padding=1)
5 self.conv2 = $n$.Conv2d(16, 32, kernel_size=3, stride=1, padding=1)
6 self.lstm = $n$.LSTM(32 * 8 * 8, hidden_size, num_layers, batch_first=True)
7 self.fc = $n$.Linear(hidden_size, num_classes)
@@
12 x = $n$.functional.relu(self.conv1(x))
13 x = $n$.functional.max_pool2d(x, 2)
14 x = $n$.functional.relu(self.conv2(x))
15 x = $n$.functional.max_pool2d(x, 2)
@@ 14.2.4 Conclusion
Radar systems, particularly modern FMCW radars, generate point cloud data that captures 3D spatial information. Point clouds are sets of points in a 3D space that represent the surfaces and objects detected by the radar. Unlike traditional image-based data, point clouds are unordered and irregular, making them challenging to process using conventional deep learning models. PointNet [QiPoi17] and PointNet++ are two architectures specifically designed to handle point cloud data, making them well-suited for tasks such as object detection, segmentation, and classification in radar systems.
@@ 15.1.1 PointNet Architecture
5class PointNet($n$.Module):
@@
8 self.fc1 = $n$.Linear(3, 64)
9 self.fc2 = $n$.Linear(64, 128)
10 self.fc3 = $n$.Linear(128, 1024)
11 self.fc4 = $n$.Linear(1024, 512)
12 self.fc5 = $n$.Linear(512, 256)
13 self.fc6 = $n$.Linear(256, num_classes)
@@ 15.2.1 Hierarchical Learning in PointNet++
1class PointNetPlusPlus($n$.Module):
@@
4 self.fc1 = $n$.Linear(3, 64)
5 self.fc2 = $n$.Linear(64, 128)
6 self.fc3 = $n$.Linear(128, 1024)
7 self.fc4 = $n$.Linear(1024, 512)
8 self.fc5 = $n$.Linear(512, 256)
9 self.fc6 = $n$.Linear(256, num_classes)
@@ 16.1.1 Self-Attention Mechanism
At the core of the Transformer architecture is the self-attention mechanism [zhou2019r], which allows the model to examine the entire input sequence simultaneously and assign different attention weights to each element in the sequence. This enables the model to focus on important parts of the data, regardless of their position in the sequence.
@@
5class SelfAttention($n$.Module):
@@
10 self.values = $n$.Linear(embed_size, embed_size, bias=False)
11 self.keys = $n$.Linear(embed_size, embed_size, bias=False)
12 self.queries = $n$.Linear(embed_size, embed_size, bias=False)
13 self.fc_out = $n$.Linear(embed_size, embed_size)
@@ 16.2.1 Transformer-based Models for Radar Data
1class TransformerRadarModel($n$.Module):
@@
4 self.self_attention_layers = $n$.ModuleList([
@@ 5 SelfAttention(embed_size, heads) for _ in range(num_layers)
7 self.fc = $n$.Linear(embed_size, num_classes)
@@ 16.3.1 Self-Attention in 3D Point Cloud Data
5class TransformerPointCloud($n$.Module):
@@
8 self.attention = $n$.MultiheadAttention(embed_dim=embed_size, num_heads=heads)
9 self.fc1 = $n$.Linear(embed_size, 512)
10 self.fc2 = $n$.Linear(512, num_classes)
@@ 16.4.2 Transformers for Long-Range Dependencies
1class PointNetPlusPlusTransformer($n$.Module):
@@
5 self.attention = $n$.MultiheadAttention(embed_dim=embed_size, num_heads=heads)
6 self.fc = $n$.Linear(embed_size, num_classes)
@@ 17.1.2 ResNet for Feature Extraction
ResNet (Residual Networks) is one of the most popular pre-trained models for feature extraction [almabdy2021feature]. It introduces skip connections, allowing the network to bypass certain layers and avoid the vanishing gradient problem. This architecture enables the model to be very deep, making it effective for capturing intricate patterns in radar data.
@@
9resnet = $n$.Sequential(*list(resnet.children())[:-1])